OpenStudy (anonymous):

the polar equation r=6cosθ - 2sinθ is the polar form of a circle. determine the center and radius of this circle algebraically

OpenStudy (tkhunny):

Have you considered multiplying by "r"?

OpenStudy (anonymous):

No I haven't. what will that do?

OpenStudy (tkhunny):

It's close to magic. Give it a go.

OpenStudy (anonymous):

Would you get \[r ^{2}=6\cos \Theta r - 2 \sin \Theta r?\] Is there more you can do?

OpenStudy (tkhunny):

It might help if you would write it this way: \(r^{2} = 6(r\cos(\theta)) - 2(r\sin(\theta))\)

OpenStudy (anonymous):

Okay. When I look at it this way do the 6 and 2 happen to be the center? R^2 I know in normal circles equals the radius squared does this have significance here

OpenStudy (tkhunny):

So close!!! Not quite. Change it to rectangular coordinates and check out your theory.

OpenStudy (anonymous):

Dang! How do I change them to regular coordinates?

OpenStudy (tkhunny):

\(r^{2} = x^{2} + y^{2}\) \(x = r\cos(\theta)\) \(y = r\sin(\theta)\)

OpenStudy (anonymous):

So would it be \[6=r \cos \Theta and -2=r \sin \theta \]

OpenStudy (tkhunny):

?? Something fell apart, there \(r^{2} = 6r\cos(\theta) - 2r\sin(\theta) \rightarrow x^{2} + y^{2} = 6x - 2y\) Give that a good hard look until you see it.

OpenStudy (anonymous):

Hmm.. Is that a formula? It isn't ringing a bell? :( I know it should be something

OpenStudy (tkhunny):

It was ONLY the three substitutions I gave you. \(r^{2} = 6r\cos(\theta) - 2r\sin(\theta)\) Given \(r^{2} = x^{2} + y^{2}\), we have \(x^{2} + y^{2} = 6r\cos(\theta) - 2r\sin(\theta)\) Given \(x = r\cos(\theta)\), we have \(x^{2} + y^{2} = 6x - 2r\sin(\theta)\) Given \(y = r\sin(\theta)\), we have \(x^{2} + y^{2} = 6x - 2y\) Substitution. That's all it is.

OpenStudy (anonymous):

Oh okay. I see it now with the break down. Thank you. How does the center and radius come out of that. I think I am seeing how to get the center but not the radius

OpenStudy (tkhunny):

Time for your best "Completing the Square" skills. How's your algebra? Now's the time to prove it!

OpenStudy (anonymous):

I think they are pretty good. I will work it out and then show what I got

OpenStudy (anonymous):

Okay. So I got\[x ^{2}- 6x +9 or (x-3)^{2}. \]. For Y I got\[-y ^{2}-2y+1\]which isn't a perfect square. Does the y remain positive?

OpenStudy (tkhunny):

You DO have to do them together. \(x^{2} + y^{2} = 6x - 2y\) Set it up \(x^{2} - 6x + y^{2} + 2y = 0\) Complete the Square on x \((x^{2} - 6x + 9) + y^{2} + 2y = 9\) Complete the Square on y \((x^{2} - 6x + 9) + (y^{2} + 2y + 1) = 9 + 1\) Rewrite \((x-3)^{2} + (y+1)^{2} = 10\) You seem to have the idea. You just have to be a little more organized.

OpenStudy (anonymous):

Okay. Perfect. So this should mean looking at the equation that the center is is at (3,-1). Would the radius be 10 or root 10? The equation looks so familiar.

OpenStudy (tkhunny):

Rewrite: \((x-3)^{2} + (y+1)^{2} = \left(\sqrt{10}\right)^{2}\) What we need to learn is what the original equation says. This will save us from converting the world to Rectangular Coordinates. \(r=6cosθ - 2sinθ\) See how that +6 turns into -3? See how that -2 turns into +1? See how 3 and 1 turn into 10? \(3^{2} + 1^{2}\) These are the lessons we must learn.

OpenStudy (anonymous):

Aww okay. It is beginning to make sense. They are squaring those numbers. I can start to see it. Math is so tricky yet cool.